The water vapor is pressurized at first 140 kPa and its volume is 150 liters, then compressed so that the volume becomes 30 liters following the law pV1,2 = C. determine the work and final pressure
1
Expert's answer
2020-10-13T18:02:34-0400

Initial pressure(p1)=140x103 Pa

Initial Volume (V1)=150 Lit = 0.150 m3

Final volume(V2)= 30 Lit=0.030 m3

Final pressure(p2)=?

Gas undergoes a polytropic process given by pV1.2 =constant

Therefore, for process 1-2 we can write,

"p_1 \times {V_1}^{1.2} = p_2 \times {V_2}^{1.2}"

"{P_2}=\frac { p_1 \times {V_1}^{1.2} }{{V_2}^{1.2}}"

"{P_2}=\frac { 140 \times {0.150}^{1.2} }{{0.030}^{1.2}}" =965.81 kPa

The work done during poly tropic process is given by

"W=\frac {p_1 \times V_1-p_2 \times V_2}{n-1}"

"W=\frac {(140\times (0.150)-(965.81)\times (0.030)}{1.2-1}" =-38.8715 kJ (-ve sign indicates work done on the gas)

Work done on the gas is W=39.8715 kJ